# On the inclusion of non-concurrent controls in platform trials with a futility interim analysis

### v1

@@ Appendix A Derivation of the bias of the model-based treatment effect estimator for Arm 2
Let $\bar{y}_{ks}$ denote the sample mean in arm $k$ and period $s$. Assume there is a stepwise time trend in the platform trial, where the strength of the trend is parameterized by $\lambda$, i.e., the mean response increases by $\lambda$ at the onset of period 2. Hence, the sample means in period 1 $\bar{y}_{01}$ and $\bar{y}_{11}$ follow normal distributions with means $\mu_{0}$ and $\mu_{1}$, respectively. In period 2, the sample means $\bar{y}_{02}$, $\bar{y}_{12}$, and $\bar{y}_{22}$ follow normal distributions with means $\mu_{0}+\lambda$, $\mu_{1}+\lambda$, and $\mu_{2}+\lambda$, respectively.
Denote by $Z_{11}=(\bar{y}_{11}-\bar{y}_{01})/\sigma\sqrt{1/n_{11}+1/n_{01}}$ the Z-statistic from the interim analysis of Arm 1. The arm is dropped for futility if the p-value from the z-test is larger than a futility boundary $\alpha_{1}$, hence if $Z_{11}<\Phi^{-1}(1-\alpha_{1})=c_{1}$. The probability of Arm 1 to be stopped at the interim analysis is $\text{I\kern-1.38742ptP}(Z_{11}<c_{1})=\Phi(c_{1}-\theta_{1}/(\sigma\sqrt{1/n_{11}+1/n_{01}}))$. To simplify the notation, define $\gamma=c_{1}-\theta_{1}/(\sigma\sqrt{1/n_{11}+1/n_{01}})$.
@@
Based on the result of the interim analysis, the estimator $\tilde{\theta}_{2}$ is either given by the separate analysis (if $Z_{11}<c_{1}$), or by the regression model (if $Z_{11}\geq c_{1}$). Let $I(\cdot)$ denote the indicator function. The expected value $E\left[\tilde{\theta}_{2}\right]$ is then given by:
$E\left[\tilde{\theta}_{2}\right]=E\left[\left[\underbrace{\bar{y}_{22}-\bar{y}_{02}}_{(1)}\right]\cdot I\{Z_{11}<c_{1}\}\right]+E\left[\left[\underbrace{\bar{y}_{22}-(1-\varrho)\bar{y}_{02}}_{(2)}-\underbrace{\varrho(\bar{y}_{01}+\bar{y}_{12}-\bar{y}_{11})}_{(3)}\right]\cdot I\{Z_{11}\geq c_{1}\}\right]$
@@
$\text{(i)}\ E[(1)\cdot I\{Z_{11}<c_{1}\}]=((\mu_{2}+\lambda)-(\mu_{0}+\lambda))\cdot\Phi(\gamma)$
$\text{(ii)}\ E[(2)\cdot I\{Z_{11}\geq c_{1}\}]=(\mu_{2}+\lambda)(1-\Phi(\gamma))-(1-\varrho)(\mu_{0}+\lambda)(1-\Phi(\gamma))$
$\text{(iii)}\ E[(3)\cdot I\{Z_{11}\geq c_{1}\}]=-\varrho(\mu_{1}+\lambda)(1-\Phi(\gamma))+\varrho\mu_{1}(1-\Phi(\gamma))-\varrho\mu_{0}(1-\Phi(\gamma))+\varrho\cdot\text{Cov}\left[(\bar{y}_{11}-\bar{y}_{01}),I\{Z_{11}\geq c_{1}\}\right]$
where in (iii) we have also used $E\left[(\bar{y}_{11}-\bar{y}_{01})\cdot I\{Z_{11}\geq c_{1}\}\right]=E\left[\bar{y}_{11}-\bar{y}_{01}\right]\cdot E\left[I\{Z_{11}\geq c_{1}\}\right]+\text{Cov}\left[(\bar{y}_{11}-\bar{y}_{01}),I\{Z_{11}\geq c_{1}\}\right]$.
@@
$E\left[\tilde{\theta}_{2}\right]=\theta_{2}+\varrho\cdot\text{Cov}\left[(\bar{y}_{11}-\bar{y}_{01}),I\{Z_{11}\geq c_{1}\}\right]$
Note that a covariance between a continuous variable $Y$ and a binary variable $X$, which takes the value 1 with probability $p$, i.e., $E[X]=p$, can be written as $\text{Cov}[X,Y]=p(1-p)\cdot(E[Y\mid X=1]-E[Y\mid X=0])$. Moreover, note that $I(Z_{11}\geq c_{1})=I\left(\bar{y}_{11}-\bar{y}_{01}\geq c_{1}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\right)$. Hence, we can rewrite the covariance $\text{Cov}\left[(\bar{y}_{11}-\bar{y}_{01}),I\{Z_{11}\geq c_{1}\}\right]$ as follows:
@@
$\varrho\cdot(1-\Phi(\gamma))\cdot\Phi(\gamma)\cdot\left(E\left[\bar{y}_{11}-\bar{y}_{01}\mid\bar{y}_{11}-\bar{y}_{01}\geq c_{1}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\right]-E\left[\bar{y}_{11}-\bar{y}_{01}\mid\bar{y}_{11}-\bar{y}_{01}<c_{1}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\right]\right)$
Since $\bar{y}_{11}-\bar{y}_{01}\sim\mathcal{N}\left(\theta_{1},\sigma^{2}\left(\frac{1}{n_{01}}+\frac{1}{n_{11}}\right)\right)$, we can use the following formulae for the expectations of the truncated normal distribution: Let $X\sim N(\mu,\sigma^{2})$, then $E[X\mid X>a]=\mu+\sigma\frac{\phi((a-\mu)/\sigma)}{1-\Phi((a-\mu)/\sigma)}$ and $E[X\mid X<b]=\mu-\sigma\frac{\phi((b-\mu)/\sigma)}{\Phi((b-\mu)/\sigma)}$. Therefore, it follows:
@@
$\varrho\cdot\left(1-\Phi(\gamma)\right)\cdot\Phi(\gamma)\cdot\left(\left(\theta_{1}+\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(\frac{c_{1}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}-\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}{1-\Phi\left(\frac{c_{1}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}-\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}\right)-\left(\theta_{1}-\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(\frac{c_{1}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}-\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}{\Phi\left(\frac{c_{1}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}-\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}\right)\right)$
@@
$E\left[\tilde{\theta}_{2}-\theta_{2}\right]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\phi\left(\gamma\right)$
@@
Thus, the expectation of $\tilde{\theta}_{2}$ conditional on $Z_{11}\geq c_{1}$ is defined as follows:
$E[\tilde{\theta}_{2}\mid Z_{11}\geq c_{1}]=\frac{E\left[\left[\bar{y}_{22}-(1-\varrho)\bar{y}_{02}-\varrho(\bar{y}_{01}+\bar{y}_{12}-\bar{y}_{11})\right]\cdot I\{Z_{11}\geq c_{1}\}\right]}{\text{I\kern-1.38742ptP}(Z_{11}\geq c_{1})}$
@@
$E\left[\tilde{\theta}_{2}-\theta_{2}\mid Z_{11}\geq c_{1}\right]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(\gamma\right)}{1-\Phi\left(\gamma\right)}$
@@
$E\left[\tilde{\theta}_{2}-\theta_{2}\mid Z_{11}\geq c_{1}\right]_{\theta_{1}=0}=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(c_{1}\right)}{\alpha_{1}}$

### v2

@@ Appendix A Derivation of the bias of the model-based treatment effect estimator for Arm 2
Let $\bar{y}_{ks}$ denote the sample mean in arm $k$ and period $s$. Assume patient $j$ enters the trial at time $t_{j}$, and is enrolled in period $s_{j}$. Further, assume a time trend defined by a function $f(t_{j})$ of the patient entry time $t_{j}$ ($j=1,\ldots,N$), which is the same for all arms and additive on the model scale, as described in Bofill Roig et al. [3]. For each period $s$, define $m_{s}:=E[f(t_{j})\mid s_{j}=s]$ and assume that $E[f(t_{j})\mid k_{j}=k,s_{j}=s]=m_{s}$ for all arms $k$ that are open for randomization in period $s$. We additionally assume that the variances of the outcomes $y_{j}$ are equal across periods. This approximately holds when the strength of the time trend is small relative to the response variance conditional on time. Hence, the sample means in period 1 $\bar{y}_{01}$ and $\bar{y}_{11}$ follow normal distributions with means $\mu_{0}+m_{1}$ and $\mu_{1}+m_{1}$, respectively. In period 2, the sample means $\bar{y}_{02}$, $\bar{y}_{12}$, and $\bar{y}_{22}$ follow normal distributions with means $\mu_{0}+m_{2}$, $\mu_{1}+m_{2}$, and $\mu_{2}+m_{2}$, respectively.
Denote by $Z_{11}=(\bar{y}_{11}-\bar{y}_{01})/\sigma\sqrt{1/n_{11}+1/n_{01}}$ the Z-statistic from the interim analysis of Arm 1. The arm is dropped for futility if the p-value from the z-test is larger than a futility boundary $\alpha_{F}$ ($0\leq\alpha_{F}\leq 1$), hence if $Z_{11}<\Phi^{-1}(1-\alpha_{F})=c_{F}$. Equivalently, the arm is dropped for efficacy if the p-value from the z-test is smaller than an efficacy boundary $\alpha_{E}$ ($0\leq\alpha_{E}\leq 1$, $\alpha_{E}<\alpha_{F}$), hence if $Z_{11}>\Phi^{-1}(1-\alpha_{E})=c_{E}$.
The probability of Arm 1 to be stopped at the interim analysis is $\text{I\kern-1.38742ptP}(Z_{11}<c_{F})+\text{I\kern-1.38742ptP}(Z_{11}>c_{E})=\Phi(c_{F}-\theta_{1}/(\sigma\sqrt{1/n_{11}+1/n_{01}}))+\Phi(\theta_{1}/(\sigma\sqrt{1/n_{11}+1/n_{01}})-c_{E})=\text{I\kern-1.38742ptP}(stop)$.
@@
Based on the result of the interim analysis, the estimator $\tilde{\theta}_{2}$ is either given by the separate analysis (if $Z_{11}<c_{F}$ or $Z_{11}>c_{E}$), or by the regression model (if $c_{F}\leq Z_{11}\leq c_{E}$). Let $I(\cdot)$ denote the indicator function. The expected value $E\left[\tilde{\theta}_{2}\right]$ is then given by:
$E\left[\tilde{\theta}_{2}\right]=E\left[\left[\underbrace{\bar{y}_{22}-\bar{y}_{02}}_{(1)}\right]\cdot I\{Z_{11}<c_{F}\text{ or }Z_{11}>c_{E}\}\right]+E\left[\left[\underbrace{\bar{y}_{22}-(1-\varrho)\bar{y}_{02}}_{(2)}-\underbrace{\varrho(\bar{y}_{01}+\bar{y}_{12}-\bar{y}_{11})}_{(3)}\right]\cdot I\{c_{F}\leq Z_{11}\leq c_{E}\}\right]$
@@
$\text{(i)}\ E[(1)\cdot I\{Z_{11}<c_{F}\text{ or }Z_{11}>c_{E}\}]=((\mu_{2}+m_{2})-(\mu_{0}+m_{2}))\cdot\text{I\kern-1.38742ptP}(stop)$
$\text{(ii)}\ E[(2)\cdot I\{c_{F}\leq Z_{11}\leq c_{E}\}]=(\mu_{2}+m_{2})(1-\text{I\kern-1.38742ptP}(stop))-(1-\varrho)(\mu_{0}+m_{2})(1-\text{I\kern-1.38742ptP}(stop))$
$\text{(iii)}\ E[(3)\cdot I\{c_{F}\leq Z_{11}\leq c_{E}\}]=-\varrho(\mu_{1}+m_{2})(1-\text{I\kern-1.38742ptP}(stop))+\varrho(\mu_{1}+m_{1})(1-\text{I\kern-1.38742ptP}(stop))-\varrho(\mu_{0}+m_{1})(1-\text{I\kern-1.38742ptP}(stop))+$
$\varrho\cdot\text{Cov}\left[(\bar{y}_{11}-\bar{y}_{01}),I\{c_{F}\leq Z_{11}\leq c_{E}\}\right]$
where in (iii) we have also used $E\left[(\bar{y}_{11}-\bar{y}_{01})\cdot I\{c_{F}\leq Z_{11}\leq c_{E}\}\right]=E\left[\bar{y}_{11}-\bar{y}_{01}\right]\cdot E\left[I\{c_{F}\leq Z_{11}\leq c_{E}\}\right]+\text{Cov}\left[(\bar{y}_{11}-\bar{y}_{01}),I\{c_{F}\leq Z_{11}\leq c_{E}\}\right]$.
@@
$E\left[\tilde{\theta}_{2}\right]=\theta_{2}+\varrho\cdot\text{Cov}\left[(\bar{y}_{11}-\bar{y}_{01}),I\{c_{F}\leq Z_{11}\leq c_{E}\}\right]$
Note that $I(c_{F}\leq Z_{11}\leq c_{E})=I\left(c_{F}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\leq\bar{y}_{11}-\bar{y}_{01}\leq c_{E}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\right)$. Moreover, note that the covariance $\text{Cov}\left[X,I(a\leq X\leq b)\right]=\text{I\kern-1.38742ptP}(a\leq X\leq b)(E[X|a\leq X\leq b]-E[X])$.
Hence, it follows:
@@
$\varrho\cdot(1-\text{I\kern-1.38742ptP}(stop))\left(E\left[\bar{y}_{11}-\bar{y}_{01}\mid c_{F}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\leq\bar{y}_{11}-\bar{y}_{01}\leq c_{E}\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\right]-\theta_{1}\right)$
Since $\bar{y}_{11}-\bar{y}_{01}\sim\mathcal{N}\left(\theta_{1},\sigma^{2}\left(\frac{1}{n_{01}}+\frac{1}{n_{11}}\right)\right)$, we can use the following formula for the expectation of the truncated normal distribution: Let $X\sim N(\mu,\sigma^{2})$, then $E[X\mid a<X<b]=\mu-\sigma\frac{\phi((b-\mu)/\sigma)-\phi((a-\mu)/\sigma)}{\Phi((b-\mu)/\sigma)-\Phi((a-\mu)/\sigma)}$. Therefore, it follows:
@@
$\varrho\cdot\left(1-\text{I\kern-1.38742ptP}(stop)\right)\left(\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(c_{F}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)-\phi\left(c_{E}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}{\Phi\left(c_{E}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)-\Phi\left(c_{F}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}\right)$
@@
$E\left[\tilde{\theta}_{2}-\theta_{2}\right]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\left(\phi\left(c_{F}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)-\phi\left(c_{E}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)\right)$
@@
Thus, the expectation of $\tilde{\theta}_{2}$ conditional on $c_{F}\leq Z_{11}\leq c_{E}$ is defined as follows:
$E[\tilde{\theta}_{2}\mid c_{F}\leq Z_{11}\leq c_{E}]=\frac{E\left[\left[\bar{y}_{22}-(1-\varrho)\bar{y}_{02}-\varrho(\bar{y}_{01}+\bar{y}_{12}-\bar{y}_{11})\right]\cdot I\{c_{F}\leq Z_{11}\leq c_{E}\}\right]}{\text{I\kern-1.38742ptP}(c_{F}\leq Z_{11}\leq c_{E})}$
@@
$E\left[\tilde{\theta}_{2}-\theta_{2}\mid c_{F}\leq Z_{11}\leq c_{E}\right]=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(c_{F}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)-\phi\left(c_{E}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}{\Phi\left(c_{E}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)-\Phi\left(c_{F}-\frac{\theta_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}\right)}$
@@
$E\left[\tilde{\theta}_{2}-\theta_{2}\mid c_{F}\leq Z_{11}\leq c_{E}\right]_{\theta_{1}=0}=\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\frac{\phi\left(c_{F}\right)-\phi\left(c_{E}\right)}{\alpha_{F}-\alpha_{E}}$